1+cose=

1+cose=

Grade:10

5 Answers

Abhishek rana
8 years ago
sec
 
KARTHIK
8 years ago
1+cose=sec
OUTPASS
8 years ago
sec
SHANMUKESHWAR
8 years ago
1+1/sec
=sec+1/sec
=tan/sec            sec^2-tan^2=1  tan^2=1+sec^2
=sin/cos/sec      1/cos=sec 
=sin/sec/sec
sin.sec/sec=sin
 
SHANMUKESHWAR
8 years ago
sorry sine